OpenStudy (madisonkayelee):

What is the slope of the following curve at the point (2, 20)? f(x)=x^4+3x^2-8 A. 10 B. 18 C. 32 D. 44

OpenStudy (3mar):

May I help?

OpenStudy (madisonkayelee):

yea

OpenStudy (jtresatti2):

its 3

OpenStudy (3mar):

Thank you.

OpenStudy (3mar):

you mean C?

OpenStudy (madisonkayelee):

no dont listen to him

OpenStudy (3mar):

Sorry I thught that is you so What did you pick?

OpenStudy (madisonkayelee):

nothing yet im lost

OpenStudy (3mar):

It is easy In Sha' Allah

OpenStudy (3mar):

the slope of this function...can you calculate it?

OpenStudy (madisonkayelee):

do i plug 2 in for x and solve

OpenStudy (3mar):

no man you should firstly find out the slope and then plug x=2 in the result

OpenStudy (3mar):

@madisonkayelee

OpenStudy (madisonkayelee):

i have no calculator

OpenStudy (3mar):

Just do it and I will calculate it for you

OpenStudy (madisonkayelee):

−3√5

OpenStudy (3mar):

\[f(x)=x^4+3x^2-8\] \[f'(x)=4x^3+6x\] now you can plug x=2 in, you get \[f'(2)=4(2)^3+6(2)=??\]

OpenStudy (3mar):

Thank you for the medal! What matters is that you get the idea and can do solving such this yourself.